OpenStudy (anonymous):

create an equation that has an x-intercept of 1/2 and a y-intercept of -3

OpenStudy (anonymous):

You can find the gradient using those two coordinates - an x-intercept of 1/2 means the point (1/2, 0) lies on the line, and a y-intercept of -3 means the point (0, -3) lies on the line. Use rise/step (the difference in y coordinates divided by the difference in x coordinates) to find the gradient. Make that your 'm'. The y-intercept is given, so make -3 your 'c'. Substitute your newfound 'm' and 'c' into the general equation of the line: y = mx + c

OpenStudy (anonymous):

thank you but I really need a more specific answer

OpenStudy (anonymous):

Okay, so to find the gradient, you'll take the coordinates (1/2, 0) and (0, -3). Rise/Step means the gradient is (-3-0) / (0 - 1/2) = (-3) / (-1/2) = 6. Let m = 6. The y intercept is -3, so let c = -3. Therefore the equation of the line is y = 6x - 3. To test, substitute in x = 1/2 and you'll see that y = 6(1/2) - 3 = 0, which is the answer we wanted (because that gives our coordinate (1/2, 0) <- the x-intercept. Hope that helps.

OpenStudy (anonymous):

when the x and y intercepts of a straight line are given then the equation of the line is given by:- (x/a)+(y/b)=1, where a and b are x and y intercepts respectively

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

ur welcome